OpenStudy (johnnydicamillo):

Find the Derivative of 7e^(x+3)+e^-1

OpenStudy (johnnydicamillo):

\[7e ^{x+3}+e ^{-1}\]

OpenStudy (ageta):

what is the x

OpenStudy (ageta):

7(2.718281828459045x+3)+2.718281828459045−1

OpenStudy (johnnydicamillo):

what do you mean? I need to find the derivative.

OpenStudy (ageta):

if you are Evaluate i need to find the x

OpenStudy (johnnydicamillo):

No, we do not need to find x just f'(x)

OpenStudy (ageta):

sorry i don't know

OpenStudy (loser66):

e^-1 is constant, then derivative =0

OpenStudy (loser66):

\(7e^{x+3}= 7(e^x.e^3)\)

OpenStudy (ageta):

ok

OpenStudy (loser66):

so that it is 7e^3 e^x, the first 2 term are constant , so let them as they are., just the last one need to take derivative

OpenStudy (loser66):

(e^x)' = e^x

OpenStudy (loser66):

so that the whole thing is \(7e^{x+3}\)